OpenStudy (anonymous):

Medals! Pizza Palace offers five sizes: personal (8in), small (10in) medium (12in), large (14in), and extra large (16in). the cost C(d), of making a pepperoni pizza can be expressed as a function of the diameter, d, of the pizza. d C 8 2.51 10 3.93 12 5.65 14 7.70 16 10.05 what is the average rate of change in cost from an 8 in to a 16 in pizza

OpenStudy (anonymous):

(8, 2.51) (16, 10.05) Average Rate of Change is the same thing as slope, so we can use the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{10.05-2.51}{16-8}\) Subtract 10.05 - 2.51, what do you get?

OpenStudy (anonymous):

@AdamLamebert

OpenStudy (anonymous):

7.54?

OpenStudy (anonymous):

Yes, now subtract 16 - 8

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Yes, so we have \(\dfrac{7.54}{8}\) Now divide 7.54 / 8

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

0.94

OpenStudy (anonymous):

Yeah, that's the average rate of change.

OpenStudy (anonymous):

HI